40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 16 12:58:50 2014, manasa, Summary, IOO, Ringdown recap  Ringdown_815.pdfMCringdown_cum.pdfcum_plot.png
    Reply  Mon Jun 16 15:32:12 2014, rana, Summary, IOO, Ringdown recap  
Message ID: 10042     Entry time: Mon Jun 16 12:58:50 2014     Reply to this: 10043
Author: manasa 
Type: Summary 
Category: IOO 
Subject: Ringdown recap  

A recap of the ringdown measurements made at the 40m in mid 2012, the hardware that was installed for the same and results from the measurements.

IMC ringdown:

Hardware installed 
A trans PD was installed (elog 7122).
This PD does not exist in the trans path anymore.

Measurements
The polarity in the MC servo was flipped with the MC WFS disabled and the PD trans signal was used to look at the cavity ringdown. 
Ringdown time = 13 us
Cavity finesse from the measurement = 453 (inconsistent with actual finesse). 

Attachment: Ringdown measurement and fits

PMC ringdown:

Hardware installed
The AOM was installed before the PMC. The AOM was driven by the driver installed on the PSL table. (RF power ~1.5W @ 80MHz @1.0V modulation input). An RF switch was installed to control the AOM driver input. ZASWA-2-50DR+ was installed.
Note: The AOM was used by the ISS crew after this. So the RF switch has been removed and the AOM is no more a part of the ringdown setup.

Measurements
No measurements were made as we could not obtain relevant TTL signals to control the switch remotely.

Attachment 1: Ringdown_815.pdf  13 kB  | Hide | Hide all
Ringdown_815.pdf
Attachment 2: MCringdown_cum.pdf  212 kB  Uploaded Mon Jun 16 14:00:08 2014  | Hide | Hide all
MCringdown_cum.pdf
Attachment 3: cum_plot.png  40 kB  Uploaded Mon Jun 16 14:00:28 2014  | Hide | Hide all
cum_plot.png
ELOG V3.1.3-